# Quantile Reward Policy Optimization: Alignment with Pointwise Regression and Exact Partition Functions

### v1

@@ Appendix I Noise amplitude in the transformed reward estimate
L.6 Sampling for reference completions and evaluations
@@
When sampling evaluation completions, we use the recommended temperature and top-$p$ for each model, that is a temperature of 0.6 and top-$p$ of 0.9 for Llama 8B Tülu 3 SFT and a temperature of 0.7 and top-$p$ 0.9 for Mistral 7B Instruct v0.2. The parameters of Llama 8B Tülu 3 SFT are taken from its default generation configuration file and the parameters of Mistral 7B Instruct v0.2 are a recommended choice within API servers.777 The data was available on https://openrouter.ai/, but the website changed interface and does not show this data anymore. These are also the parameters suggested when asking Le Chat.
All the sampling is done with vLLM (Kwon et al., 2023) and we keep all the other sampling parameters as default. In early experiments we generated up to 200 completions per prompt, and since this is an embarrassingly parallel problem, our codebase includes a scalable sampling pipeline with data parallelism to effectively scale this task to a cluster scale.
To generate completions for the LeetCode dataset, we modify the original prompt to make the model aware of the initialization code that runs before its suggested solution. This code includes generic imports and definitions of helper classes (see Figures 9, 10, and 11). The imports and definitions are used in the template solution code that the model gets in its prompt and in the test suite, and are therefore critical in its crafting of the problem solution. The original initialization code is not constant across problems and we found it to lack robustness and to be subject to bad practices such as having several wildcard import statements. We therefore combined and refactored all initialization prefixes into a single unified one shown in Figures 10 and 11.
@@
L.7 Distribution shift settings
It is well known that although policy fitting methods can be used to fit the optimal policy with any data distribution, they do exhibit different levels of performance depending on the distribution shift between the model being trained and the completions in the training dataset. For example, it can be recommended to perform SFT on the chosen completions when the completions in the preference dataset are not from the model being aligned. We therefore let the distribution shift setting be a hyperparameter and select 6 different distribution shift settings to put all algorithms in their best setting. We consider the SFT-chosen and no-SFT cases where the model is first fine-tuned on the chosen completions or not (our initial models are already instruction fine-tuned, this is an extra fine-tuning to reduce distribution shift). Subsequent to these two settings we have the offline setting (data from the dataset) and two off-policy settings, where we replace the chosen and rejected completions in the dataset (keeping the prompts) with the completions generated by the model before training (thus $2\times(1+2)=6$ settings). We refer to the off-policy settings by (i) off-policy-best when we follow Xu et al. (2024b) in picking the chosen (resp. rejected) from the best (resp. worst) out of 6 completions scored by the ArmoRM reward model, and by (ii) off-policy-random when we only pick them by comparing and annotating 2 random completions from the 6 completions. Unlike popular belief, we observe that DPO obtains better results on off-policy-random than on off-policy-best in multiple cases (see Table 12 for example).
@@
L.8 Hardware and software acceleration
We train all the models (SFT and RL fine-tuning) with the same distributed training setup. We distribute training with HuggingFace Accelerate (Gugger et al., 2022) using the DeepSpeed (Aminabadi et al., 2022) plugin at ZeRO stage 1 (optimizer state partitioning) over 8 GPUs from 2 nodes with 4 NVIDIA GH200 each.
Our SFT trainer and RL fine-tuning trainer (same trainer supporting DPO, SimPO, REBEL, and QRPO) are based on the HuggingFace TRL (von Werra et al., 2020) SFT and DPO trainers.
L.9 Baseline algorithms and hyperparameters
@@
For SFT (when fine-tuning on the chosen completions) we use a learning rate of $75e-$ 7 and train for one epoch in the chat task and three epochs in the code task (smaller dataset) The SFT loss is computed on all the tokens in the sequence (prompt and completion). This is the default in the TRL trainer. We apply gradient clipping by norm with value $10.0$ (to prevent potential spikes, though the norm stabilizes below $10.0$ very early in training).
@@
For RL fine-tuning, we sweep over three learning rates $[1e-7,3e-7,1e-6]$ and train for one epoch. We run three $\beta$ values for each algorithm, following a $\log$ scale, unless specific hyperparameters are recommended by the authors of the respective algorithms. For the chat task, for DPO we sweep over $[0.01,0.03,0.1]$ which is the same magnitude as in Meng et al. (2024) and are also commonly used in the literature, for REBEL we sweep over $[1e-6,1e-4,1e-2]$ as done by Gao et al. (2024) in a similar experimental setting using UltraFeedback and ArmoRM, for SimPO we sweep over $[2,2.5,10]$ as recommended by Meng et al. (2024) keeping the margin at 0.5 as recommended by their public codebase to maintain the same budget of hyperparameters for all methods, and for QRPO we sweep over $[3e-4,0.001,0.003]$ which we found to be a good range in preliminary experiments on another chat dataset. The code task is harder and we expect a worse offline-to-online generalization requiring to aim for a more conservative optimal policy and change in the rewards. Therefore we adjust the range of $\beta$ values used in the sweep for each method depending on the impact of $\beta$ on the method to obtain similar impacts: for REBEL we increase $\beta$ by 2 orders of magnitude and we sweep over $[1e-4,1e-2,1.0]$, one order to be more conservative and one order to match the difference between rewards which increases from the chat task by around one order of magnitude since REBEL is sensitive to the reward scale to preserve the signal-to-regularization ratio (higher $\beta$ values have also been used by Gao et al. (2024) in their summarization experiment); For QRPO we increase $\beta$ by only one order of magnitude and sweep over $[0.003,0.01,0.03]$, we only need one order to be more conservative and do not need to take into account the change in reward scale as the quantile reward is invariant to it; For DPO we keep the same values for $\beta$, since this parameter is not tied to an external reward scale (model learns an internal reward implicitly based on preferences). Moreover, we do not adjust for additional conservatism, as $\beta=1$ is too conservative. Finally for SimPO, the training completions are at least twice shorter therefore $\beta$ is already effectively increased in the loss and the values do not follow a log scale so we can’t increase by an order of magnitude and thus we keep the same values.
@@
To estimate the percentile rewards for QRPO, we use $n=1$ reference model rewards for Magpie-Air, $n=3$ for UltraFeedback, and $n=20$ for LeetCode. Note that when using $n=1$, the quantile reward is obtained by comparing the reward of the training sample to the reward of the single reference reward and yields a binary quantile reward of 00 or $1$.
@@
For downstream evaluations, on the chat task we further evaluate the best checkpoint of every algorithm on AlpacaEval 2.0 (Li et al., 2023; Dubois et al., 2024) and report the length-controlled win rate against GPT-4 preview (gpt-4-1106-preview denoted as gpt4_turbo in the AlpacaEval repository) as the final performance of an algorithm on a (model, dataset) configuration. AlpacaEval 2.0 is a widely used benchmark by the community, assessing broad conversational skills across a wide range of queries with a total of 805 questions. On the code task, the reward on the test split of the evaluation set can already be considered as downstream as they measure the performance of the model on the popular LeetCode coding task.
L.11 Statistical significance
@@
L.12 Common implementation pitfalls and experiment protocol limitations
@@
  • There was a bug in SFT Trainer of the TRL library (von Werra et al., 2020) where when the padding token was set to be the same as the EOS token as shown in their examples, the EOS token would be masked out in the loss like the padding token, leading to models never predicting and end of sequence during the SFT and then at inference time generating non-ending sequences. This bug has since been fixed in TRL (#3200, #3328).
  • Popular models diverge in the EOS used for pre-training and post-training with their HuggingFace config remaining with the pre-training EOS and leading them to generate non-ending sequences out of the box. (E.g., the Magpie-Align models trained with the Axolotl library)(EOS and chat template).
  • An extra EOS token is added to the sequence after applying the chat template if it already contains an EOS in the DPO implementation of TRL. This bug is still present in TRL at the time of writing. (Added EOS after chat template.)
@@
  • In the HuggingFace Transformers library (Wolf et al., 2020) there was a bug leading to the gradient norm to scale with the number of accumulation steps, which is solved at the time of writing (#35808).
@@
Some benchmarks or models do not specify the sampling parameters for their evaluations. For example, AlpacaEval (Li et al., 2023) does not specify a decoding configuration, resulting in different works reporting performance on AlpacaEval with different decoding strategies even when using the same base model making the comparison across works very hard. For example, for the same Llama 3 8B Xu et al. (2024b) use greedy decoding with a repetition penalty on AlpacaEval while Meng et al. (2024) perform a grid search over multiple sampling temperatures and report their results with a temperature of 0.9.
@@
We believe that new methods should compare to baselines while maintaining the same hyperparameter budget. This is because if a new method introduces a new hyperparameter that increases the computational budget to find its optimum, baselines could as well find better optima by searching over more hyperparameters. However, we find that this is not always followed. For example, while Meng et al. (2024) perform a comprehensive hyperparameter search for their baselines, they still use four times more hyperparameters for SimPO than for DPO.
@@
In offline alignment, at least with the current algorithms, there is no accurate metric to predict online performance on downstream evaluations. For example, we have observed that the DPO loss, margins, or accuracy are not reliable metrics to predict online performance. This has also been observed by Rafailov et al. (2024) for other policy fitting algorithms. Furthermore, the best hyperparameters cannot be selected based on the online performance on the downstream evaluations, as this practice would be subject to a selection bias. Therefore, it is important for offline alignment work to report their hyperparameter selection protocol. Yet, we have found that several research papers do not document their hyperparameter selection protocol. This also leads to reproducibility issues and discrepancies between papers that seemingly use the same experimental protocol (datasets, hyperparameters, etc.), but still report different numbers.
Appendix M Detailed empirical results
@@
Figure 12: Scaling curves in Figure 3 for different learning rates. For a fixed $\beta$ and learning rate, we still select the best checkpoint according to a validation set and report the performance on a test set, as described in our experimental protocol in Appendix L.
Table 12: Llama 8B Tülu 3 SFT, Magpie-Air dataset. The best models selected across distribution shifts are marked with * according to rewards and LC-rewards as described in the selection protocol in Appendix L. We underline the values determining the selection shortlist using the validation split (valid).
@@
offpolicy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1734}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1761}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1671}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1706}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ 0.01 3e-7 320
offpolicy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1774}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1804}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1682}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0009}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1716}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ 0.01 1e-6 480

### v2

@@ Appendix I Noise amplitude in the partition function estimate for the initial reward
M.6 Sampling for reference completions and evaluations
@@
When sampling evaluation completions, we use the recommended temperature and top-$p$ for each model, that is a temperature of 0.6 and top-$p$ of 0.9 for Llama 8B Tülu 3 SFT and a temperature of 0.7 and top-$p$ 0.9 for Mistral 7B Instruct v0.2. The parameters of Llama 8B Tülu 3 SFT are taken from its default generation configuration file and the parameters of Mistral 7B Instruct v0.2 are a recommended choice within API servers. [^7] The data was available on https://openrouter.ai/, but the website changed interface and does not show this data anymore. These are also the parameters suggested when asking Le Chat.
All the sampling is done with vLLM (kwon2023efficient) and we keep all the other sampling parameters as default. In early experiments we generated up to 200 completions per prompt, and since this is an embarrassingly parallel problem, our codebase includes a scalable sampling pipeline with data parallelism to effectively scale this task to a cluster scale.
To generate completions for the LeetCode dataset, we modify the original prompt to make the model aware of the initialization code that runs before its suggested solution. This code includes generic imports and definitions of helper classes (see Figures 11, 12, and 13). The imports and definitions are used in the template solution code that the model gets in its prompt and in the test suite, and are therefore critical in its crafting of the problem solution. The original initialization code is not constant across problems and we found it to lack robustness and to be subject to bad practices such as having several wildcard import statements. We therefore combined and refactored all initialization prefixes into a single unified one shown in Figures 12 and 13.
@@
M.7 Distribution shift settings
It is well known that although policy fitting methods can be used to fit the optimal policy with any data distribution, they do exhibit different levels of performance depending on the distribution shift between the model being trained and the completions in the training dataset. For example, it can be recommended to perform SFT on the chosen completions when the completions in the preference dataset are not from the model being aligned. We therefore let the distribution shift setting be a hyperparameter and select 6 different distribution shift settings to put all algorithms in their best setting. We consider the SFT-chosen and no-SFT cases where the model is first fine-tuned on the chosen completions or not (our initial models are already instruction fine-tuned, this is an extra fine-tuning to reduce distribution shift). Subsequent to these two settings we have the offline setting (data from the dataset) and two off-policy settings, where we replace the chosen and rejected completions in the dataset (keeping the prompts) with the completions generated by the model before training (thus $2\times(1+2)=6$ settings). We refer to the off-policy settings by (i) off-policy-best when we follow xu_magpie_2024 in picking the chosen (resp. rejected) from the best (resp. worst) out of 6 completions scored by the ArmoRM reward model, and by (ii) off-policy-random when we only pick them by comparing and annotating 2 random completions from the 6 completions. Unlike popular belief, we observe that DPO obtains better results on off-policy-random than on off-policy-best in multiple cases (see Table 12 for example).
@@
M.8 Hardware and software acceleration
We train all the models (SFT and RL fine-tuning) with the same distributed training setup. We distribute training with HuggingFace Accelerate (accelerate) using the DeepSpeed (deepspeed) plugin at ZeRO stage 1 (optimizer state partitioning) over 8 GPUs from 2 nodes with 4 NVIDIA GH200 each.
Our SFT trainer and RL fine-tuning trainer (same trainer supporting DPO, SimPO, REBEL, and QRPO) are based on the HuggingFace TRL (vonwerra2022trl) SFT and DPO trainers.
M.9 Baseline algorithms and hyperparameters
@@
For SFT (when fine-tuning on the chosen completions) we use a learning rate of 5e75e-7 and train for one epoch in the chat task and three epochs in the code task (smaller dataset) The SFT loss is computed on all the tokens in the sequence (prompt and completion). This is the default in the TRL trainer. We apply gradient clipping by norm with value $10.0$ (to prevent potential spikes, though the norm stabilizes below $10.0$ very early in training).
@@
For RL fine-tuning, we sweep over three learning rates [1e7,3e7,1e6][1e-7,3e-7,1e-6] and train for one epoch. We run three $\beta$ values for each algorithm, following a $\log$ scale, unless specific hyperparameters are recommended by the authors of the respective algorithms. For the chat task, for DPO we sweep over $[0.01,0.03,0.1]$ which is the same magnitude as in meng2024simpo and are also commonly used in the literature, for REBEL we sweep over [1e6,1e4,1e2][1e-6,1e-4,1e-2] as done by gao_rebel_2024 in a similar experimental setting using UltraFeedback and ArmoRM, for SimPO we sweep over $[2,2.5,10]$ as recommended by meng2024simpo keeping the margin at 0.5 as recommended by their public codebase to maintain the same budget of hyperparameters for all methods, and for QRPO we sweep over [3e4,0.001,0.003][3e-4,0.001,0.003] which we found to be a good range in preliminary experiments on another chat dataset. The code task is harder and we expect a worse offline-to-online generalization requiring to aim for a more conservative optimal policy and change in the rewards. Therefore we adjust the range of $\beta$ values used in the sweep for each method depending on the impact of $\beta$ on the method to obtain similar impacts: for REBEL we increase $\beta$ by 2 orders of magnitude and we sweep over [1e4,1e2,1.0][1e-4,1e-2,1.0], one order to be more conservative and one order to match the difference between rewards which increases from the chat task by around one order of magnitude since REBEL is sensitive to the reward scale to preserve the signal-to-regularization ratio (higher $\beta$ values have also been used by gao_rebel_2024 in their summarization experiment); For QRPO we increase $\beta$ by only one order of magnitude and sweep over $[0.003,0.01,0.03]$, we only need one order to be more conservative and do not need to take into account the change in reward scale as the quantile reward is invariant to it; For DPO we keep the same values for $\beta$, since this parameter is not tied to an external reward scale (model learns an internal reward implicitly based on preferences). Moreover, we do not adjust for additional conservatism, as $\beta=1$ is too conservative. Finally for SimPO, the training completions are at least twice shorter therefore $\beta$ is already effectively increased in the loss and the values do not follow a log scale so we can’t increase by an order of magnitude and thus we keep the same values.
@@
To estimate the percentile rewards for QRPO, we use $n=1$ reference model rewards for Magpie-Air, $n=3$ for UltraFeedback, and $n=20$ for LeetCode. Note that when using $n=1$, the quantile reward is obtained by comparing the reward of the training sample to the reward of the single reference reward and yields a binary quantile reward of 0 or $1$.
@@
For downstream evaluations, on the chat task we further evaluate the best checkpoint of every algorithm on AlpacaEval 2.0 (alpaca_eval; dubois2024length) and report the length-controlled win rate against GPT-4 preview (gpt-4-1106-preview denoted as gpt4_turbo in the AlpacaEval repository) as the final performance of an algorithm on a (model, dataset) configuration. AlpacaEval 2.0 is a widely used benchmark by the community, assessing broad conversational skills across a wide range of queries with a total of 805 questions. On the code task, the reward on the test split of the evaluation set can already be considered as downstream as they measure the performance of the model on the popular LeetCode coding task.
M.11 Statistical significance
@@
M.12 Common implementation pitfalls and experiment protocol limitations
@@
  • There was a bug in SFT Trainer of the TRL library (vonwerra2022trl) where when the padding token was set to be the same as the EOS token as shown in their examples, the EOS token would be masked out in the loss like the padding token, leading to models never predicting and end of sequence during the SFT and then at inference time generating non-ending sequences. This bug has since been fixed in TRL (#3200, #3328).
  • Popular models diverge in the EOS used for pre-training and post-training with their HuggingFace config remaining with the pre-training EOS and leading them to generate non-ending sequences out of the box. (E.g., the Magpie-Align models trained with the Axolotl library)(Links to EOS and chat template).
  • An extra EOS token is added to the sequence after applying the chat template if it already contains an EOS in the DPO implementation of TRL. This bug is still present in TRL at the time of writing. (Links to Added EOS after chat template.)
@@
  • In the HuggingFace Transformers library (wolf-etal-2020-transformers) there was a bug leading to the gradient norm to scale with the number of accumulation steps, which is solved at the time of writing (#35808).
@@
Some benchmarks or models do not specify the sampling parameters for their evaluations. For example, AlpacaEval (alpaca_eval) does not specify a decoding configuration, resulting in different works reporting performance on AlpacaEval with different decoding strategies even when using the same base model making the comparison across works very hard. For example, for the same Llama 3 8B xu_magpie_2024 use greedy decoding with a repetition penalty on AlpacaEval while meng2024simpo perform a grid search over multiple sampling temperatures and report their results with a temperature of 0.9.
@@
We believe that new methods should compare to baselines while maintaining the same hyperparameter budget. This is because if a new method introduces a new hyperparameter that increases the computational budget to find its optimum, baselines could as well find better optima by searching over more hyperparameters. However, we find that this is not always followed. For example, while meng2024simpo perform a comprehensive hyperparameter search for their baselines, they still use four times more hyperparameters for SimPO than for DPO.
@@
In offline alignment, at least with the current algorithms, there is no accurate metric to predict online performance on downstream evaluations. For example, we have observed that the DPO loss, margins, or accuracy are not reliable metrics to predict online performance. This has also been observed by rafailov2024scaling for other policy fitting algorithms. Furthermore, the best hyperparameters cannot be selected based on the online performance on the downstream evaluations, as this practice would be subject to a selection bias. Therefore, it is important for offline alignment work to report their hyperparameter selection protocol. Yet, we have found that several research papers do not document their hyperparameter selection protocol. This also leads to reproducibility issues and discrepancies between papers that seemingly use the same experimental protocol (datasets, hyperparameters, etc.), but still report different numbers.
Appendix N Detailed empirical results
@@
Figure 14: Scaling curves in Figure 3 for different learning rates. For a fixed $\beta$ and learning rate, we still select the best checkpoint according to a validation set and report the performance on a test set, as described in our experimental protocol in Appendix M.
Table 12: Llama 8B Tülu 3 SFT, Magpie-Air dataset. The best models selected across distribution shifts are marked with * according to rewards and LC-rewards as described in the selection protocol in Appendix M. We underline the values determining the selection shortlist using the validation split (valid).
@@
off-policy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1734}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1761}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1671}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1706}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ 0.01 3e-7 320
off-policy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1774}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1804}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1682}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0009}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1716}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ 0.01 1e-6 480